### v1

# Lipschitz regularity of definable sets

Abstract.
In this paper, we prove that for any $X\subset\mathbb{R}^{n}$ that is a $d$-dimensional definable set in an o-minimal structure and is Lipschitz regular at any $p\in X$, the following items are equivalent:
  1. (1)
    $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$;
  2. (2)
    $C_{4}(X,p)=C_{3}(X,p)$ for any $p\in X$;
  3. (3)
    $C_{4}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$;
  4. (4)
    $C_{5}(X,p)$ is a $d-$dimensional linear subspace for any $p\in X$;
  5. (5)
    $X$ is $C^{1}$ smooth.
Key words and phrases:
Lipschitz geometry, analytic sets, regularity
2010 Mathematics Subject Classification:
14B05; 32S50
The third named author was partially supported by CNPq-Brazil grant 310438/2021-7. This work was supported by the Serrapilheira Institute (grant number Serra – R-2110-39576).
1. Introduction
A subject of interest for many mathematicians is to know how simple is the topology of complex analytic sets. For example, if topological regularity implies analytic regularity. In general this does not occur, but Mumford in [13] showed a result in this direction, the result can be formulated as follows: a topological regular complex surface in $\mathbb{C}^{3}$, with isolated singularity, is smooth. The second author in his thesis [15] (see also [16] and [4]) proved a version of the Mumford’s Theorem, he showed that if a complex analytic set is Lipschitz regular then it is smooth. Thus, if we think in the real case, for each positive integer $d$ we have the following question.
  1. Question LR($d$) Let $X\subset\mathbb{R}^{n}$ be a real analytic set with dimension $d$. Suppose that $X$ is Lipschitz regular at $0\in X$. Is it true that $X$ is $C^{1}$ smooth at 00?
In the paper [17], it was proved that above question has a positive answer if and only if $d=1$. In particular, it was proved that for $d>1$, $X=\{(x_{1},...,x_{d+1})\in\mathbb{R}^{d+1};\,x_{3}^{3}=x_{1}^{3}+x_{2}^{3}\}$ is a bi-Lipschitz homeomorphic to $\mathbb{R}^{d}$.
Thus, in order to get $C^{1}$ smoothness of real analytic sets, we have to ask more than Lipschitz regularity.
In this article, we restrict ourselves to the problem of characterizing $C^{1}$ submanifolds in the context of o-minimal structures in terms of their tangent cones defined below:
Definition 1.1.
Let $X\subset\mathbb{R}^{n}$ be a set such that $p\in\overline{X}$. We define $C_{i}(X,p)$, $i=3,4$ and $5$, as follows:
  1. (1)
    $v\in C_{3}(X,p)$ if there are a sequence of points $\{x_{i}\}_{i}\subset X$ tending to $p$ and a sequence of positive real numbers $\{t_{i}\}_{i}$ such that
    $\lim\limits_{i\to\infty}\frac{1}{t_{i}}(x_{i}-p)=v.$
  2. (2)
    $v\in C_{4}(X,p)$ if there are a sequence of points $\{x_{i}\}\subset{\rm Reg}_{1}X$ tending to $p$ and a sequence of $\{v_{i}\}_{i}\subset$ tending to $v$ such that $v_{i}\in T_{x_{i}}X$ for all $i$. Here, ${\rm Reg}_{1}X$ denotes the set of points $p\in X$ where $X$ is a $C^{1}$ submanifold around $p$.
  3. (3)
    $v\in C_{5}(X,p)$ if there are sequences of points $\{x_{i}\}_{i},\{y_{i}\}_{i}\subset X$ tending to $p$ and a sequence of positive real numbers $\{t_{i}\}_{i}$ such that
    $\lim\limits_{i\to\infty}\frac{1}{t_{i}}(x_{i}-y_{i})=v.$
Characterizing $C^{1}$ submanifolds of $\mathbb{R}^{n}$ in terms of their third and fifty tangent cones has already been studied by many authors, see, for example, [2, 10, 11, 12, 19], and the survey [1]. For instance, it was proved in [12, Theorem 4.7] the following very interesting characterization:
Theorem 1.2.
Let $X\subset\mathbb{R}^{n}$ be a locally closed definable set in an o-minimal structure. Then, $X$ is $C^{1}$ smooth if and only if $C_{5}(X,p)=C_{3}(X,p)$ for any $p\in X$.
In this article, we prove the following results.
Theorem 3.4.
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional definable set in an o-minimal structure. Assume that $X$ is Lipschitz regular at any $p\in X$. Then the following items are equivalent.
  1. (1)
    $C_{4}(X,p)=C_{3}(X,p)$ for any $p\in X$;
  2. (2)
    $C_{4}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$;
  3. (3)
    $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$;
  4. (4)
    $X$ is $C^{1}$ smooth.
Theorem 3.2.
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional locally closed definable set in an o-minimal structure. Then $X$ is $C^{1}$ smooth if and only if $X$ is LNE at any $p\in X$, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and continuously depends on $p$.
As a consequence, we obtain [12, Theorem 4.7], which was stated above as Theorem 1.2 (see Corollary 3.10).
In the parametrized case, Nuño-Ballesteros and Mendes in [14] proved that for a germ of a real analytic mapping $F\colon(\mathbb{R}^{n},0)\to(\mathbb{R}^{k},0)$ that is a lipeomorphism, its image $X=Im(F)$ is smooth in 00.
So, we have the following natural question:
Question 1.
Let $F\colon(\mathbb{R}^{n},0)\to(\mathbb{R}^{k},0)$ be the germ of a real analytic mapping and $X=Im(F)$. If $X$ is Lipschitz regular at 00, is $X$ $C^{1}$ smooth at 00?
Unfortunately, this question has a negative answer as given in the following example.
Example 1.3.
Let $F\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be the mapping given by $F(x,y)=(x,y^{3},(x+y)^{3})$. Then $X=Im(F)$ is Lipschitz regular at 00, but $X$ is not smooth at 00.
The details explaining why $X=Im(F)$ is Lipschitz regular at 00 and is not smooth at 00 are presented in Proposition 3.19.
Acknowledgements. We would like to thank Guillermo Sanchis Peñafort for helping us find the example of Proposition 3.19. We would also like to thank Alexandre Fernandes, Euripedes C. da Silva, Juan José Nuño Ballesteros and Lev Birbrair.
2. Preliminaries
Lemma 2.1.
Let $X\subset\mathbb{R}^{n}$ be a definable set such that $p\in\overline{X}$. Then $C_{3}(X,p)\subset C_{4}(X,p)\subset C_{5}(X,p)$.
Proof.
First observe that $C_{3}(X,p)=C_{3}(Reg_{1}(X),p)$. Now given $v\in C_{3}(X,p)=C_{3}(Reg_{1}(X),p)$ we have that there is a curve $\alpha:[0,\epsilon)\rightarrow Reg_{1}(X)$ such that, among other things, $\alpha^{\prime}(0)=v$, but since $\alpha(t)\in Reg_{1}(X)$, we have that $\alpha^{\prime}(t)\in T_{\alpha(t)}X$. As $\alpha\in C^{1}$, then $\lim_{t\to 0^{+}}\alpha^{\prime}(t)=\alpha^{\prime}(0)=v$. Take $v\in C_{4}(X,p)$, since $v_{k}\in T_{x_{k}}X$, then there is a curve $\alpha:(-\epsilon,\epsilon)\to X$ with $\alpha(0)=x_{k}$, such that $v_{k}=\alpha^{\prime}(0)$. But this means that there are sequences $(y_{k_{i}})_{i}$ and $(z_{k_{i}})_{i}$ in $(-\epsilon,\epsilon)$ with $(y_{k_{i}})_{i}\leq 0\leq(z_{k_{i}})_{i}$ and $\lim_{i\to\infty}(y_{k_{i}})_{i}=0=\lim_{i\to\infty}(z_{k_{i}})_{i}$, so that
$v_{k}=\alpha^{\prime}(0)=\lim_{i\to\infty}{\alpha((z_{k_{i}})_{i})-\alpha((y_{k_{i}})_{i})\over(z_{k_{i}})_{i}-(y_{k_{i}})_{i}},$
then for all $v_{k}$, fixed $i>>1$ we get $a_{k}=(z_{k_{i}})_{i}$ and $b_{k}=(y_{k_{i}})_{i}$ so that
$\|v_{k}-{\alpha(a_{k})-\alpha(b_{k})\over a_{k}-b_{k}}\|<{1\over 2i},$
then $v\in C_{5}(X,p)$. ∎
Lemma 2.2.
Let $X\subset\mathbb{R}^{n}$ be a pure $d$-dimensional definable set such that $p\in\overline{X}$. Then $C_{4}(X,p)$ is the union of all $d$-dimensional linear subspaces $T$ such that there is a sequence of points $\{x_{i}\}_{i}\subset{\rm Reg}_{1}X$ tending to $p$ with $\lim\limits_{i\to\infty}T_{x_{i}}X=T$.
Proof.
Let $T$ be a $d$-dimensional linear subspaces such that there is a sequence of points $\{x_{j}\}_{j}\subset{\rm Reg}_{1}X$ tending to $p$ with $\lim\limits_{j\to\infty}T_{x_{j}}X=T$. We may assume that $\delta(T_{x_{j}}X,T)<{1\over j}$, where $\delta$ is the usual Grassmannian metric. Let $\{w_{1},w_{2},\cdots,w_{d}\}$ be a basis of $T$, which, we may assume satisfies $\|w_{i}\|=1$ for $1\leq i\leq k$. Since $\delta(T_{x_{j}}X,T)<\frac{1}{j}$, there exists $v_{i_{j}}\in T_{x_{j}}X$ such that $\|v_{i_{j}}-w_{i}\|<\frac{1}{i}$. Thus, given $w\in T$, we have $w=a_{1}w_{1}+\cdots+a_{d}w_{d}$ for constants $a_{1},\cdots,a_{d}\in\mathbb{R}$, and therefore the sequence of vectors $w_{j}=a_{1}v_{1_{j}}+\cdots+a_{n}v_{k_{j}}\in T_{x_{j}}X$ converges to $w$. Thus, $w\in C_{4}(X,p)$.
Now take a vector $v\in C_{4}(X,p)$, then there is a sequence $v_{i}\to v$, as in the definition, we have that the sequence $(T_{x_{i}}X)$ converges to a d-dimensional subspace of $\mathbb{R}^{n}$, say $S$. So short of normalizing the vectors of $(v_{i})_{i}$ and $v$, we have to
${1\over i}>\delta(S,T_{x_{i}}X)\geq sup_{\|w\|=1,w\in T_{x_{i}}X}\{\|w-\pi_{S}(w)\|\}\geq\|v_{i}-\pi_{S}(v_{i})\|,$
which proves the result.
We finish this subsection by recalling the definition of bi-Lipschitz homeomorphism and the invariance of the tangent cone under bi-Lipschitz homeomorphisms obtained in the paper [16].
Definition 2.3.
We say that a mapping $f\colon X\subset\mathbb{R}^{n}\to Y\subset\mathbb{R}^{m}$ is a bi-Lipschitz homeomorphism if $f$ is a bijection and there exists a constant $K\geq 1$ such that
$\frac{1}{K}\|x-y\|\leq\|f(x)-f(y)\|\leq K\|x-y\|\quad\forall x,y\in X.$
Theorem 2.4 (Sampaio [16], Theorem 3.2).
Let $X,Y\subset\mathbb{R}^{n+1}$ be two germs of analytic subsets. If $\varphi\colon(\mathbb{R}^{n+1},X,0)\to(\mathbb{R}^{n+1},Y,0)$ is a bi-Lipschitz homeomorphism, then there is a bi-Lipschitz homeomorphism $d\varphi\colon(\mathbb{R}^{n+1},C_{3}(X,0),0)\to(\mathbb{R}^{n+1},C_{3}(Y,0),0)$.
3. Lipschitz regularity of definable sets
Definition 3.1.
A subset $X\subset\mathbb{R}^{n}$ is called Lipschitz regular at $p\in X$ if there is an open neighborhood $U$ of $p$ in $X$ which is bi-Lipschitz homeomorphic to an Euclidean ball.
3.1. Lispchitz regularity and the third tangent cone
Theorem 3.2.
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional locally closed definable set in an o-minimal structure. Then $X$ is $C^{1}$ smooth if and only if $X$ is LNE at any $p\in X$, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$.
Proof.
Clearly, if $X$ is $C^{1}$ smooth, then $X$ is LNE at any $p\in X$, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$.
Reciprocally, we assume that $X$ is LNE at any $p\in X$, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$. Fix $p\in X$. We may assume that $p=0$.
Let $\Pi\colon\mathbb{R}^{n}\to C_{3}(X,0)$ be the orthogonal projection onto $C_{3}(X,0)$. We choose linear coordinates $(x,y)$ in $\mathbb{R}^{n}$ such that $C_{3}(X,0)=\{(x,y)\in\mathbb{R}^{n};y=0\}\cong\mathbb{R}^{d}$. With this identification, $\pi(x,y)=x$. Let $L={\rm Ker}(\Pi)=\{(x,y)\in\mathbb{R}^{n};x=0\}$.
By hypotheses, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $q\in X$, and depends continuously on $q$. Thus, we can choose a small enough open neighbourhood $U$ of 00 such that $C_{3}(X,q)\cap L=\{0\}$ for all $q\in\overline{U}\cap X$. In particular, $\Pi|_{C_{3}(X,q)}\colon C_{3}(X,q)\to C_{3}(X,0)$ is a linear isomorphism.
Since $X$ is a locally closed set, we can shrink $U$, if necessary, so that $X\cap\overline{U}$ is a compact subset.
By shrinking $U$ again, if necessary, we may assume that $(X\cap U)\cap P^{-1}(\pi(q))$ is a finite subset for all $q\in X\cap U$. Indeed, if this was not possible, there would be $q\in X\cap U$ and a sequence $\{q_{j}\}_{j}\subset(X\setminus\{\tilde{q}\})\cap\Pi^{-1}(\Pi(q))$ such that $\lim q_{j}=\tilde{q}\in X\cap\overline{U}$. Taking a subsequence, if necessary, we may assume that $\lim\frac{q_{j}-\tilde{q}}{\|q_{j}-\tilde{q}\|}=v$. Then $v\in C_{3}(X,\tilde{q})\cap L$, which is a contradiction, since $C_{3}(X,\tilde{q})\cap L=\{0\}$ and $v\not=0$.
From above, we obtain, in particular, that $(X\cap U)\cap L$ is a finite subset and, thus by shrinking $U$ again, if necessary, we may assume that $(X\cap\overline{U})\cap L=\{0\}$.
Claim 3.2.1.
$\pi=\Pi|_{A}$ is an open mapping, where $A=X\cap U$.
Proof of Claim 3.2.1.
The proof of this claim shares its structure with the proof of Lemma 3.3 in [10] (see also [12, Lemma 4.1]).
Fix $q\in A$. Since $X$ is locally closed, there is a $r>0$ small enough such that $V:=\overline{B_{r}^{n}(q)}\cap X\subset A$ is a compact subset, and the topological boundary $\partial V$ of $V$ in $X$ lies on $\partial B_{r}^{n}(q)$, which implies that $q\not\in\partial V.$ Moreover, for a small enough $r>0$, we can assume that
$\pi(q)\not\in\pi(\partial V)$
otherwise, there exists a sequence of positive numbers $\{r_{k}\}$ tending to 00 such that for each $k$, there is a point $q_{k}\in X\cap\partial\overline{B_{r_{k}}^{n}(q)}$ with $\pi(q_{k})=\pi(q)$. So, extracting a subsequence if necessary, we can assume that $\lim\limits_{k\to\infty}\frac{q_{k}-q}{\|q_{k}-q\|}=v\not=0$. This implies that $v\in L\cap C_{3}(X,q)$, which is a contradiction, since $C_{3}(X,q)\cap L=\{0\}$. Since $\pi(\partial V)$ is a compact set, there is $s>0$ such that
(1) $\overline{B_{s}^{d}(\pi(q))}\cap\pi(\partial V)=\emptyset.$
It is enough to show that $\pi(q)$ is an interior point of $\pi(V)$ in $P$.
Thus, suppose by contradiction that $\pi(q)$ is not an interior point of $\pi(V)$. Then ${B_{\delta}^{d}(\pi(q))}\not\subset\pi(V)$, for any $\delta>0$. In particular, there is a point
$x\in{B_{s/2}^{d}(\pi(q))}\setminus\pi(V).$
Since $x\not\in\pi(V)$ and $\pi(V)$ is compact, for $t=dist(x,\pi(V))$, we have that $\overline{B_{t}^{d}(x)}\subset C_{3}(X,0)$ intersects $\pi(V)$ while $B_{t}^{d}(x)\cap\pi(V)=\emptyset$. Moreover, since $q\in V$, we have $t\leq\|x-\pi(q)\|<s/2$, and if $x^{\prime}\in\overline{B_{t}^{d}(x)}$ then
$\|x^{\prime}-\pi(q)\|\leq\|x^{\prime}-x\|+\|x^{\prime}-\pi(q)\|\leq t+s/2<s,$
which yields
$\overline{B_{t}^{d}(x)}\subset B_{s}^{d}(\pi(q)).$
Thus $\overline{B_{t}^{d}(x)}\cap\pi(\partial V)=\emptyset$ by (1). Take $\tilde{x}\in\overline{B_{t}^{d}(x)}\cap\pi(V)$ and $\tilde{p}\in\pi^{-1}(\tilde{x})\cap V$. Note that $\tilde{p}\not\in\partial V$, so $\tilde{p}$ is an interior point of $V$, and hence $C_{3}(V,\tilde{p})=C_{3}(X,\tilde{p})$.
Since $B_{t}^{d}(x)\cap\pi(V)=\emptyset$, no point of $V$ is contained in the cylinder $C:=\pi^{-1}(B_{t}^{d}(x))$. This implies that $\ell\subset T_{\tilde{p}}\partial C$ for each line $\ell\subset C_{3}(X,\tilde{p})$ passing through the origin. In fact, let $\ell\subset C_{3}(X,\tilde{p})$ be a line passing through the origin, then we have two arcs $\gamma_{1},\gamma_{2}\colon[0,\varepsilon)\to V$ such that $\gamma_{i}(\tau)-\tilde{p}=(-1)^{i}\tau u+o(\tau)$ for $i=1,2$, where $u\in\mathbb{R}^{n}\setminus\{0\}$ satisfies =$\ell=\{\tau u;\tau\in\mathbb{R}\}$. Thus, suppose that $\ell\not\subset T_{\tilde{p}}\partial C$, then the line $\tilde{p}+\ell=\{\tilde{p}+\tau u;\tau\in\mathbb{R}\}$ and $C$ intersect transversally at $\tilde{p}$, this forces the image of $\gamma_{1}$ or $\gamma_{2}$ to intersect $C$, which is a contradiction, since $V\cap C=\emptyset$. Therefore, $\ell\subset T_{\tilde{p}}\partial C$. Since $C_{3}(X,\tilde{p})$ is linear, we obtain that $C_{3}(X,\tilde{p})\subset T_{\tilde{p}}\partial C$.
Since $L\subset T_{\tilde{p}}\partial C$ and $C_{3}(X,\tilde{p})\oplus L=\mathbb{R}^{n}$, we obtain that $T_{\tilde{p}}\partial C=\mathbb{R}^{n}$, which is a contradiction, since $T_{\tilde{p}}\partial C$ is a $(n-1)$-dimensional linear subspace.
Therefore, $\pi(q)$ is an interior point of $\pi(V)$ and this finishes the proof of Claim 3.2.1.
Claim 3.2.2.
$\pi$ is an injective mapping.
Proof of Claim 3.2.2.
Let $k=\max\{\#(\pi^{-1}(x));x\in\pi(A)\}$ and $E=\{x\in\pi(A);k=\#(\pi^{-1}(x))\}$. We have that $E$ is an open subset of $\pi(A)$ and $\pi\colon\pi^{-1}(E)\to E$ is a cover mapping. Since #$\#(\pi^{-1}(0))=1$, we have that $k>1$ if and only if $\pi(A)\setminus E\not=\emptyset$.
Assume that $k>1$ and let $x\in E$. Let $z\in\pi(A)\setminus E$ such that $r={\rm dist}(x,\pi(A)\setminus E)=\|x-z\|$. Then, $B_{r}(x)\subset E$ and there are $k$ Lipschitz mappings $f_{1},...,f_{k}\colon B_{r}(x)\to\mathbb{R}^{n-d}$ such that $Graph(f_{i})\subset A$, $i=1,...,k$. We have that each $f_{i}$ extends to a Lipschitz mapping $\tilde{f}_{i}\colon\overline{B_{r}(x)}\to\mathbb{R}^{n-d}$. Since $z\not\in E$, then #$\#(\pi^{-1}(z))<k$. Thus, there are different $i_{1},i_{2}\in\{1,...,k\}$ such that $\tilde{f}_{i_{1}}(z)=\tilde{f}_{i_{2}}(z)$. After reordering the indices, if necessary, we assume that $i_{1}=1$ and $i_{2}=2$.
Since $Graph(\tilde{f_{i}})$ is bi-Lipschitz homeomorphic to the ball $\overline{B_{r}(x)}$ (which has a half-space as its tangent cone), since $C_{3}(X,p)$ is a linear subspace at every point $p\in X$, we have that the tangent cone of $Graph(\tilde{f_{i}})$ at $\tilde{z}=(z,f(z))$ is also a half-space and is given by the graph of the function $C_{3}(\overline{B_{r}(x)},z)\ni v\mapsto(v,df(x)v)$, where $df$ represents the lateral derivative. Take two points $(y,f_{1}(y))\in Graph(\tilde{f_{1}})$ and $(y,f_{2}(y))\in Graph(\tilde{f_{2}})$ and curves $(\gamma_{i}(t))\subset Graph(\tilde{f_{i}})$, with $\gamma_{i}(0)=\tilde{z}$ for $i=1,2$. Since $C_{3}(Graph(\tilde{f_{1}}),\tilde{z})=C_{3}(Graph(\tilde{f_{2}}),\tilde{z})$, such curves can be taken so as to have the same tangent vector. But this is a contradiction, since ${|\gamma_{1}(t)-\gamma_{2}(t)|\over t}\geq C>0$.
Therefore, $k=1$ and $\pi$ is an injective mapping.
Let $\phi\colon(\mathbb{R}^{d},0)\to\mathbb{R}^{n-d}$ be the mapping-germ such that $\pi(x,\phi(x))=x$ for all small enough $x\in\mathbb{R}^{d}$.
Claim 3.2.3.
$\phi$ is a mapping of class $C^{1}$.
Proof of Claim 3.2.3.
This claim was already proved in [12, Lemma 4.2], but for completeness, we present that proof here.
Let us begin by showing that $\phi$ is continuous. Let $U$ be the domain of $\phi$, and let $a\in U$. Take a sequence $\{x_{k}\}_{k}\to a$, then by Claim 3.2.1, there exists a neighborhood $V$ of $(a,\phi(a))$ in $X$ such that $\pi|_{V}$ is open, where $\pi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$. Since $X$ is locally closed, we may assume, possibly by shrinking $V$, that $\overline{V}\subset X=\Gamma(\phi)$ is compact. As $a\in\pi(V)$, we have that for $k\gg 1$, $x_{k}\in\pi(V)$, so $\{(x_{k},\phi(x_{k}))\}_{k}\subset\overline{V}$. Therefore, possibly after passing to a subsequence, we may assume that this sequence converges, say, to $(a,b)$. Thus, since $\overline{V}\subset\Gamma_{\phi}$, it follows that $\phi(a)=b=\lim_{k\to\infty}\phi(x_{k})$. This shows that $\phi$ is continuous.
Now we show that $\phi$ is $C^{1}$.
Let $\sigma:\mathcal{L}(\mathbb{R}^{k},\mathbb{R}^{n-k})\ni\phi\mapsto\Gamma(\phi)\in\mathbb{G}^{k}_{n}$, that is, the map that associates a linear map $\phi$ to its graph. But since $\sigma$ is a homeomorphism onto its image, and since $C_{3}(X,p)$ is not orthogonal to $\mathbb{R}^{k}$, we have that $C_{3}(X,p)\in\sigma(\mathcal{L}(\mathbb{R}^{k},\mathbb{R}^{n-k}))$. Let $L(x)=\sigma^{-1}(C_{3}(X,(x,f(x))))$. But by construction, we have that $df(x)=L(x)$, which is a composition of continuous functions, therefore $f\in C^{1}$.
Therefore, $X$ is locally the graph of a $C^{1}$ smooth mapping, and thus $X$ is $C^{1}$ smooth. ∎
After proving this theorem, the following question naturally arises: is the above theorem still valid if, instead of requiring $C_{3}(X,p)$ as in the theorem, we require that $C_{4}(X,p)$ or $C_{5}(X,p)$ be a $d$-dimensional linear subspace and that $X$ is LNE at $p$, for all $p\in X$? It is not difficult to observe that under these conditions, $C_{4}(X,p)=C_{5}(X,p)$ varies continuously. The example below shows that this is not true.
Example 3.3.
Consider the set $X=\{(x,y)\in\mathbb{R}^{2};y^{2}-x^{3}\leq 0\}$, which is definable and satisfies $C_{4}(X,p)=\mathbb{R}^{2}=C_{5}(X,p)$ for all $p\in X$. Moreover, $X$ is LNE, but it is not $C^{1}$-smooth.
3.2. Lispchitz regularity and the forth tangent cone
Theorem 3.4.
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional definable set in an o-minimal structure. Then the following items are equivalent.
  1. (1)
    $X$ is Lipschitz regular at any $p\in X$ and $C_{4}(X,p)=C_{3}(X,p)$ for any $p\in X$;
  2. (2)
    $X$ is Lipschitz regular at any $p\in X$ and $C_{4}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$;
  3. (3)
    $X$ is $C^{1}$ smooth.
Proof.
$(1)\Rightarrow(2).$ Assume that $C_{4}(X,p)=C_{3}(X,p)$ for any $p\in X$. Fix $p\in X$. By hypotheses, $X$ is Lipschitz regular at $p$. By the bi-Lipschitz invariance of the tangent cones, $C_{3}(X,p)$ is lipeomorphic to $\mathbb{R}^{d}$ and, thus $C_{4}(X,p)$ is lipeomorphic to $\mathbb{R}^{d}$. Indeed, there is a lipeomorphism $\phi\colon\mathbb{R}^{d}\to C_{4}(X,p)$ such that $\phi(0)=0$. By definition, $C_{4}(X,p)$ is an union of $d$-dimensional linear subspaces. Let $T\subset C_{4}(X,p)$ be a $d$-dimensional linear subspace. In particular, $\phi|_{T}\colon T\to\mathbb{R}^{d}$ is an injective and continuous mapping and, thus $\phi|_{T}$ is an open mapping. In particular, $\phi(T)$ contains an open ball $B\subset\mathbb{R}^{d}$ centred at 00. This forces $C_{4}(X,p)=T$.
Therefore, $C_{4}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$.
$(2)\Rightarrow(3).$ Assume that $C_{4}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$. Fix $p\in X$. By hypotheses, $X$ is Lipschitz regular at $p$. By the bi-Lipschitz invariance of the tangent cones, there is a lipeomorphism $\phi\colon\mathbb{R}^{d}\to C_{3}(X,p)$ such that $\phi(0)=0$. In particular, the mapping $\phi\colon\mathbb{R}^{d}\to C_{4}(X,p)$ is a continuous and injective mapping and, thus an open mapping. This shows that $C_{3}(X,p)$ is an open subset of $C_{4}(X,p)$. Since $C_{3}(X,p)$ is a closed subset of $C_{4}(X,p)$ and $C_{4}(X,p)$ is a connected set, we have that $C_{3}(X,p)=C_{4}(X,p)$.
Therefore, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $x\in X$.
Claim 3.4.1.
$C_{4}(X,p)$ depends continuously on $p$.
Proof of Claim 3.4.1.
Suppose there exist sequences $(x_{k})_{k},(y_{k})_{k}\subset X$, both converging to $p$, but $\lim(C_{4}{X,x_{n}})_{n}=S\neq T=\lim(C_{4}{X,y_{n}})_{n}$. Note that we can assume that $(C_{4}{X,x_{n}})_{n}$ e $(C_{4}{X,y_{n}})_{n}$ converge, since the Grassmannian is compact. By Proposition 2.2 there exist $x^{\prime}_{n},y^{\prime}_{n}\in Reg_{1}(X)$ such that $d(C_{4}(X,x_{n}),T{x^{\prime}_{n}}X)<{1\over n}$, so $T{x^{\prime}_{n}}X\to S$ and similarly $T{y^{\prime}_{n}}X\to T$, but $T,S\subset C_{4}(X,p)$, and since $T$ and $S$ are subspaces of dimension $d$, it follows that $T=C_{4}(X,p)=S$. ∎
Therefore, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$ and, moreover, $X$ is LNE at any $p\in X$. By Theorem 3.2, $X$ is $C^{1}$ smooth. ∎
3.3. Lispchitz regularity and the fifth tangent cone
Proposition 3.5.
Let $X\subset\mathbb{R}^{n}$ be a closed, $d$-dimensional set definable in an o-minimal structure. If $dim(C_{5}(X,p))=d$, then $X$ is, in a neighborhood of $p$, the graph of a Lipschitz function defined on $\pi(U)$, for some open set $U$ of $X$ containing $p$.
Proof.
First, from the fact that $C_{5}(X,p)$ is definable in an o-minimal structure, we have that there exists a $d$-dimensional subspace $P$ such that $\pi:\mathbb{R}^{n}\rightarrow P$ satisfies $\pi^{-1}(0)\cap C_{5}(X,p)={0}$. Now observe that there exists a neighborhood $U$ of 00 in $X$ such that $\pi:X\cap U\rightarrow P$ is injective. Otherwise, given any neighborhood of 00, there would exist $x_{k},y_{k}\in U$ such that $\pi(x_{k})=\pi(y_{k})$ and, therefore, $\frac{x_{k}-y_{k}}{\|x_{k}-y_{k}\|}$ lies in the kernel of $\pi$. But, up to a subsequence, we can assume that $v=\lim_{k\to\infty}{x_{k}-y_{k}\over\|x_{k}-y_{k}\|}$ exists, hence $v\in C_{5}(X,0)$ and also lies in the kernel of $\pi$. Thus, on one hand $v=0$, and on the other hand, it has norm $1$, a contradiction. Therefore, such a neighborhood exists. Since $\pi:U\cap X\rightarrow\mathbb{R}^{d}$ is injective, we can define a function $f$ on $\pi(U\cap X)$ such that the graph of $f$ is $X\cap U$. It remains to show that $f$ is Lipschitz.
Suppose, for contradiction, that $f$ is not Lipschitz. Then for every given $k\in\mathbb{N}$, there exist sequences $(x_{k})_{k},(y_{k})_{k}\subset\pi(X\cap U)$ such that
${\|f(x_{k})-f(y_{k})\|\over\|(x_{k}-y_{k})\|}\geq k,$
but since $((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\over\|((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\|$ has norm 1, it converges, up to a subsequence. However,
${x_{k}-y_{k}\over\|((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\|}\leq{x_{k}-y_{k}\over\|(f(x_{k})-f(y_{k}))\|}\to 0,$
so the sequence $((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\over\|((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\|$ converges to a vector that lies in $\pi^{-1}(0)\cap C_{5}(X,0)=\{0\}$, but on the other hand, it has norm $1$, a contradiction. Therefore, $f$ is Lipschitz. ∎
The converse of the above proposition does not hold.
Example 3.6.
Let $X=\{(x,y)\in R^{2};y=|x|\}$. We have that $X$ is the graph of a Lipschitz function, but $C_{5}(X,0)=\{(x,y)\in R^{2};-|x|\leq y\leq|x|\}$.
In fact, we have the following more general result:
Theorem 3.7.
Let $X\subset\mathbb{R}^{n}$ be a locally closed $d$-dimensional definable set in an o-minimal structure. Then, $X$ is $C^{1}$ smooth if and only if $C_{5}(X,p)$ is a $d$-dimensional linear subspace and the restriction to $X$ of the orthogonal projection $\pi\colon\mathbb{R}^{n}\to C(X,p)$ is an open mapping for any $p\in X$.
Proof.
Since $C_{5}(X,p)$ is a $d$-dimensional linear subspace, then, up to a linear change of coordinates, we may assume that $C_{5}(X,p)=\{(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{n-d};y=0\}$. Let $\pi:\mathbb{R}^{n}\rightarrow C_{5}(X,p)$ be the canonical projection. We have that $\pi^{-1}(0)\cap C_{5}(X,p)=\{0\}$. But by the same argument used in Proposition 3.5, we have that $\pi|_{X\cap U}:X\cap U\rightarrow\pi(X\cap U)$ is bi-Lipschitz, and since the restriction to $X$ of the orthogonal projection $\pi\colon\mathbb{R}^{n}\to C_{5}(X,p)$ is an open mapping for any $p\in X$, then $\pi(X\cap U)\subset P$ is open. Hence, $X$ is Lipschitz regular at $p$. Moreover, since $X$ is $d$-dimensional, we have that $C_{4}(X,p)$ is a union of $d$-dimensional linear subspaces, but since $C_{4}(X,p)\subset C_{5}(X,p)$, it follows that $C_{4}(X,p)$ is a $d$-dimensional linear subspace. Therefore, by Theorem 3.4, the result follows. ∎
Remark 3.8.
We can prove a slightly more general result. Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional definable set in an o-minimal structure. Then, $X$ is $C^{1}$ smooth if and only if $X$ is a topological manifold and $C_{5}(X,p)$ is a $d$-dimensional subset for any $p\in X$.
Corollary 3.9.
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional definable set in an o-minimal structure. Then, $X$ is $C^{1}$ smooth if and only if $X$ is a topological manifold and $C_{5}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$.
Proof.
We already know from Proposition 3.5 that $\pi|_{X\cap U}:X\cap U\rightarrow\pi(X\cap U)$ is bi-Lipschitz. Therefore, $\pi:X\cap U\rightarrow C_{5}(X,p)$ is continuous and injective between two topological manifolds of the same dimension. It follows from the Invariance of dimension Theorem that $\pi|_{X\cap U}:X\cap U\rightarrow C_{5}(X,p)$ is an open mapping. The result then follows from Theorem 3.7. ∎
Since $C_{5}(X,p)=C_{3}(X,p)$ implies that $C_{5}(X,p)=C_{3}(X,p)$ is a $d$-dimensional linear subspace and using Claim 3.2.1, we obtain the main result in [12, Theorem 4.7]:
Corollary 3.10.
Let $X\subset\mathbb{R}^{n}$ be a locally closed $d$-dimensional definable set in an o-minimal structure. Then, $X$ is $C^{1}$ smooth if and only if $C_{5}(X,p)=C_{3}(X,p)$ for any $p\in X$.
Example 3.11.
Let $X=\{(x,y,z)\in\mathbb{R}^{3};z=-y^{2}(|x|-1)\}$. We have that for each $p\in X$, $X$ is around $p$ the graph of a Lipschitz function. Moreover, $C_{5}(X,0)$ is the plane $z=0$ and $X$ is not $C^{1}$ smooth at 00.
Summarizing, we have proved the following.
Theorem 3.12.
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional definable set in an o-minimal structure. Then the following items are equivalent.
  1. (1)
    $X$ is a topological manifold and $C_{5}(X,p)$ is a $d$-dimensional linear subspace set for any $p\in X$;
  2. (2)
    $X$ is Lipschitz regular at any $p\in X$ and $C_{4}(X,p)=C_{3}(X,p)$ for any $p\in X$;
  3. (3)
    $X$ is Lipschitz regular at any $p\in X$ and $C_{4}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$;
  4. (4)
    $X$ is locally closed and LNE at any $p\in X$, $C_{3}(X,p)$ is a $d$-dimensional linear subspace for any $p\in X$, and depends continuously on $p$;
  5. (5)
    $X$ is $C^{1}$ smooth.
3.4. Examples
In this Subsection, we present several results in order to show the sharpness of our hypotheses.
The condition that $X$ is Lipschitz regular at any $p\in X$ cannot be removed.
Example 3.13.
For $X=\{(x,y)\in\mathbb{R}^{2};xy=0\}$, we have that $C_{3}(X,p)=C_{4}(X,p)$ for all $p\in X$. But neither $C_{3}(X,0)$ or $C_{4}(X,0)$ is a linear subspace.
In fact, we cannot remove that condition even if we ask that is a topological manifold.
Example 3.14.
For $X=\{(x,y)\in\mathbb{R}^{2};x^{2}-y^{3}=0\}$, we have that $C_{4}(X,p)$ is a real line for all $p\in X$. But $C_{3}(X,0)$ is not a real line.
Example 3.15.
For $X=\{(x,y)\in\mathbb{C}^{2};x^{2}-y^{3}=0\}$, we have that $C_{4}(X,p)=C_{3}(X,p)$ is a complex line for all $p\in X$ and $X$ is a topological manifold, but $X$ is not $C^{1}$ smooth.
The condition that $C_{4}(X,p)$ is a linear subspace for all $p\in X$ cannot be weakened to “$C_{3}(X,p)$ being plane for all $p\in X$”.
Example 3.16.
Let $f\colon\mathbb{R}^{2}\to\mathbb{R}$ be the function given by
$f(x,y)=\left\{\begin{array}[]{ll}\frac{y^{6}}{x^{2}+y^{4}},&\mbox{ if }x^{2}+y^{4}\not=0,\\ 0,&\mbox{ if }x^{2}+y^{4}=0.\end{array}\right.$
Note that $f$ is a Lipschitz function but not a $C^{1}$ function. Indeed, for $x^{2}+y^{4}\not=0$, the partial derivatives $\frac{\partial f}{\partial x}(x,y)=\frac{-2xy^{6}}{(x^{2}+y^{4})^{2}}$ and $\frac{\partial f}{\partial x}(x,y)=\frac{6y^{5}}{x^{2}+y^{4}}-\frac{4y^{9}}{(x^{2}+y^{4})^{2}}$ are bounded around $(0,0)$ and $\frac{\partial f}{\partial x}$ cannot be extended as a continuous function at $(0,0)$. Thus, $X=Graph(f)$ is Lipschitz regular at 00. Since $X\setminus\{0\}$ is smooth, $C_{3}(X,p)$ is plane for all $p\in X\setminus\{0\}$. Moreover, $P=\{(x,y,z)\in\mathbb{R}^{3};z=0\}\subset C(X,0)$. Since $X$ is Lipschitz regular at 00, then $C(X,0)$ is lipeomorphic to $P$, and thus $C(X,0)=P$. Therefore, $X$ is Lipschitz regular at any $p\in X$, $C_{3}(X,p)$ is plane for all $p\in X$, but $X$ is not $C^{1}$-smooth at 00.
Example 3.17.
For each $\varepsilon>0$, let $X_{\varepsilon}=Graph(\varepsilon f)$, where $f$ is the function defined in Example 3.16. Since $f$ is Lipschitz, $\theta(X_{\varepsilon},0)<\frac{3}{2}$ for a small enough $\varepsilon>0$. However, $X_{\varepsilon}$ is not $C^{1}$ smooth at 00 and $C_{3}(X_{\varepsilon},p)$ is plane for all $p\in X_{\varepsilon}$.
We can obtain even an algebraic example.
Example 3.18.
Let $X=\{(x,y,z)\in\mathbb{R}^{3};z^{15}=x^{9}y^{7}+x^{7}y^{9}\}$. We have that for each $p\in X$, $X$ is Lipschitz regular at $p$ and $C_{3}(X,p)$ is a plane. However, $X$ is not $C^{1}$ smooth at 00.
We finish this section by proving that Question 1 has a negative answer.
Proposition 3.19.
Let $F\colon\mathbb{R}^{2}\to\mathbb{R}^{3}$ be the mapping given by $F(x,y)=(x,y^{3},(x+y)^{3})$. Then $X=Im(F)$ is an algebraic set, Lipschitz regular at any $p\in X$, $C(X,p)$ is a plane for any $p\in X$, but $X$ is not smooth at 00.
Proof.
Note that $Im(F)=V(f)$, where $f\colon\mathbb{R}^{3}\to\mathbb{R}$ is given by
$f(x,y,z)=(y-z)^{3}+x^{9}+3x^{6}y+3x^{3}y^{2}-3x^{6}z+21x^{3}yz+3x^{3}z^{2}.$
Thus $X$ is algebraic. Now note that $X\setminus\{0\}$ is smooth, so it is enough to analyze only at the origin. Consider the projection $\pi:\mathbb{R}^{3}\rightarrow P$ in the direction of the vector $\eta=(1,1,0)$, where $P=\{(x,y,z)\in\mathbb{R}^{3};x+y=0\}$. Notice that this is a bijection when restricted to $X$. Moreover, we can see that by projecting in a neighborhood of the direction $\eta=(1,1,0)$, we also obtain a bijective projection, and therefore the projection is a bi-Lipschitz homeomorphism.
To see that the cone is a linear subspace at every point, simply observe $C_{3}(X,0)=\{(x,y,z)\in\mathbb{R}^{3};y=z\}$. Finally, let us show that $X$ is not $C^{1}$. For this, note that the normal vectors to the tangent plane at points of the form $({1\over k},0)$ are vectors of the form $(0,1,0)$, but at $(0,0)$ the normal vector is $(0,1,-1)$. Therefore, the tangent cones, which are planes, do not vary continuously, and therefore $X$ is not of class $C^{1}$. ∎
3.5. Regularity of real analytic sets
Theorem 3.20.
Let $X\subset\mathbb{R}^{n}$ be a pure $d$-dimensional real analytic set. Then the following items are equivalent:
  1. (1)
    $X$ is LNE at any $p\in X$ and $C_{4}(X,p)$ is a $d$-dimensional linear subspace for all $p\in X$;
  2. (2)
    $C_{5}(X,p)$ is a $d$-dimensional linear subspace for all $p\in X$;
  3. (3)
    $X$ is $C^{1}$ smooth.
Proof.
Now we will show that $(1)\Rightarrow(3)$. In fact, we will show that $C_{3}(X,p)=C_{4}(X,p)$ for all $p\in X$. We may assume that $p=0$, and choose a system of linear coordinates $(x,y)$ in $\mathbb{R}^{n}$ such that $C_{4}(X,0)=\{(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{n-d}\mid y=0\}$. Let $\pi:\mathbb{R}^{n}\to C_{4}(X,0)\cong\mathbb{R}^{d}$ be the orthogonal projection. Since $X$ is a real analytic set, the multiplicity modulo 2 is well defined. Let $U$ be a sufficiently small neighborhood of 0 in $X$ such that there exists a subanalytic set $\sigma\subset C_{4}(X,0)$ satisfying that #$\#(\pi^{-1}(v)\cap U)\mod 2$ is constant for all $v\in C_{4}(X,0)\setminus\sigma$ sufficiently close to the origin.
Suppose that $C_{3}(X,p)\subsetneq C_{4}(X,p)$. Since $C_{3}(X,p)$ is closed, $W:=\text{int}(C_{4}(X,p)\setminus C_{3}(X,p))$ is nonempty. For $v\in W$ and for a sufficiently small neighborhood $U$ of 0 in $X$, we have $\pi^{-1}(tv)\cap U=\emptyset$ for all sufficiently small $t>0$. This shows that the multiplicity of $X$ at the origin is $0\mod 2$.
Claim 3.20.1.
$\pi(U\setminus Sing(X))$ is an open subset of $C_{4}(X,0)$.
Proof of Claim 3.20.1.
Since $C_{4}(X,q)$ varies continuously with $q$, shrinking $U$ if necessary, we may assume that $\delta(C_{4}(X,0),C_{4}(X,q))<1/3$ for all $q\in U$. In particular, for each $q\in U$, the restriction of $\pi$ to $C_{4}(X,q)$ is a linear isomorphism. Thus, $\pi(U\setminus\text{Sing}(X))$ is an open subset of $C_{4}(X,0)$. ∎
We may assume that $U=X\cap B^{n}(0,r)$ for some $r>0$. Let $B=\pi(U)$, $B^{\prime}=B\setminus\pi(\text{Sing}(X))$, $U^{\prime}=\pi^{-1}(B^{\prime})\cap U$, $N=\sup_{x\in B^{\prime}}\#\pi^{-1}_{U}(x)$, and $S=\{x\in B^{\prime}\mid\#\pi^{-1}_{U}(x)=N\}$. By shrinking r, if necessary, we assume that $0\in\overline{S}$ and $\pi^{-1}(0)\cap U=\{0\}$. Then $S$ is an open set, $\pi^{-1}_{U}(S)$ has exactly $N$ connected components, say $X_{1},\ldots,X_{N}$, and for each $i$, $X_{i}$ is the graph of a $C^{1}$ function $f_{i}:S\to\mathbb{R}^{n-d}$.
Given $z\in S$, the distance from $z$ to $B^{\prime}\setminus S$ is attained at some $z^{\prime}\in B^{\prime}\setminus S$, hence the ball $B_{r}(z)\subset S$, where $r=\text{dist}(z,B^{\prime}\setminus S)$. By the assumptions on the tangent cones, each $f_{i}$ also has bounded derivative. Let $\lambda\geq 1$ such that $\|Df_{i}\|\leq\lambda$ on $S$, for all $i\in\{1,\ldots,N\}$. Therefore, the functions $f_{i}:B_{r}(z)\rightarrow\mathbb{R}^{n-d}$ are lipschitz, so there exists a Lipschitz extension $\tilde{f}:\overline{B_{r}(z)}\rightarrow\mathbb{R}^{n-d}$.
Since the multiplicity (mod 2) at the origin is zero, we must have $N\geq 2$, that is, there exist two functions $f_{1}$ and $f_{2}$ such that $f_{1}(z^{\prime})=f_{2}(z^{\prime})$. Since the graph $G_{i}$ of each $f_{i}$ is bi-Lipschitz to the closed ball, it follows from Theorem 2.4 that the cone $C_{3}(G_{i},(z^{\prime},f_{i}(z^{\prime})))$ is bi-Lipschitz homeomorphic to the half-space $H=C_{3}(B_{r}(z))$. Moreover,
$C_{3}(G_{i},(z^{\prime},f_{i}(z^{\prime})))=\{(v,D_{+}(f_{i})_{x}v)\mid v\in H\}.$
But since $C_{4}(X,p)$ is a plane, we have $C_{4}(X,p)=\{(v,L(v))\mid v\in\mathbb{R}^{d}\}$, for some linear map $L$, hence $L(v)=D_{+}(f_{i})_{x}v$.
Now, let $\alpha$ be the line segment joining $z$ to $z^{\prime}$. Then both $(\alpha(t),f_{1}(\alpha(t)))$ and $(\alpha(t),f_{2}(\alpha(t)))$ are tangent, i.e. ${\|(\alpha(t),f_{1}(\alpha(t)))-(\alpha(t),f_{2}(\alpha(t)))\|\over t}\to 0$. On the other hand, we have
$d_{X}((\alpha(t),f_{1}(\alpha(t))),(\alpha(t),f_{2}(\alpha(t))))\geq 2\|\alpha(t)\|,$
and we may reparametrize so that $\|\alpha(t)\|=t$, which contradicts the fact that $X$ is LNE.
$(2)\Rightarrow(3)$ Since $C_{5}(X,p)$ is a $d$-dimensional linear subspace, then, up to a linear change of coordinates, we may assume that $C_{5}(X,p)=\{(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{n-d};y=0\}=P$. Let $\pi:\mathbb{R}^{n}\rightarrow P$ be the canonical projection. We have that $\pi^{-1}(0)\cap P=\{0\}$.
By the same argument used in Proposition 3.5, we have that $\pi|_{X\cap U}:X\cap U\rightarrow\pi(X\cap U)$ is bi-Lipschitz.
If we show that $\pi(X\cap U)$ contains an open set, then $X\cap U$, possibly after shrinking the open set, is Lipschitz regular. And since $C_{4}(X,p)\subset C_{5}(X,p)$, it follows that $C_{4}(X,p)$ is a $d$-dimensional linear subspace for every $p\in X$, and thus the result follows from Theorem 3.4.
Since $\pi:X\cap U\rightarrow\pi(X\cap U)$ is bi-Lipschitz, there exists a Lipschitz extension $\tilde{\pi}:\overline{X\cap U}\rightarrow\overline{\pi(X\cap U)}=\pi(\overline{X\cap U})$. But $X$ is analytic, and in particular, closed in $U$. Hence, $\overline{X\cap U}=X\cap\overline{U}$, and therefore $\pi(X\cap\overline{U})\subset P$ is closed and contains the origin.
Thus, $P\setminus\pi(X\cap\overline{U})$ is an open subset of $P$, and so $\dim(P\setminus\pi(X\cap\overline{U}))=\dim(P)>\dim(\sigma)$.
Now observe that $0\notin\overline{P\setminus\pi(X\cap\overline{U})}$. Indeed, take a sequence $(x_{j})_{j}\subset(P\setminus\pi(X\cap\overline{U}))\setminus\sigma$ with $x_{j}\to 0$. Then $\pi^{-1}(x_{j})\cap X=\emptyset$, so $m(X,0)=0\pmod{2}$, which contradicts the fact that $\pi:X\cap U\rightarrow\pi(X\cap U)$ is injective.
Therefore, $0\notin\overline{P\setminus\pi(X\cap\overline{U})}$, so $\pi(X\cap\overline{U})$ contains a neighborhood of the origin.
Remark 3.21.
We can prove a slightly more general result. Let $X\subset\mathbb{R}^{n}$ be a pure $d$-dimensional real analytic set. Then the following items are equivalent:
  1. (1)
    $X$ is LNE at any $p\in X$ and $C_{4}(X,p)$ is a $d$-dimensional linear subspace for all $p\in X$;
  2. (2)
    $C_{5}(X,p)$ is a $d$-dimensional subset for all $p\in X$;
  3. (3)
    $X$ is $C^{1}$ smooth.
4. Lipschitz regularity at infinity
Definition 4.1.
Let $X\subset\mathbb{R}^{n}$ be a set such that $p\in\overline{X}$. We define $C_{i}(X,\infty)$, $i=3,4$ and $5$, as follows:
  1. (1)
    $v\in C_{3}(X,\infty)$ if there are a sequence of points $\{x_{i}\}_{i}\subset X$ tending to $\infty$ and a sequence of positive real numbers $\{t_{i}\}_{i}$ such that
    $\lim\limits_{i\to\infty}\frac{x_{i}}{t_{i}}=v.$
  2. (2)
    $v\in C_{4}(X,\infty)$ if there are a sequence of points $\{x_{i}\}\subset{\rm Reg}_{1}X$ tending to $\infty$ and a sequence of $\{v_{i}\}_{i}\subset$ tending to $v$ such that $v_{i}\in T_{x_{i}}X$ for all $i$.
  3. (3)
    $v\in C_{5}(X,\infty)$ if there are sequences of points $\{x_{i}\}_{i},\{y_{i}\}_{i}\subset X$ tending to $\infty$ and a sequence of positive real numbers $\{t_{i}\}_{i}$ such that
    $\lim\limits_{i\to\infty}\frac{1}{t_{i}}(x_{i}-y_{i})=v.$
Remark 4.2.
Let $X\subset\mathbb{R}^{n}$ be a definable set. Then $C_{3}(X,\infty)\subset C_{4}(X,\infty)\subset C_{5}(X,\infty)$.
Definition 4.3.
A subset $X\subset\mathbb{R}^{n}$ is called Lipschitz regular at infinity if there is an compact subset $K$ such that $X\setminus K$ is bi-Lipschitz homeomorphic to $\mathbb{R}^{k}\setminus\overline{B_{1}(0)}$.
Definition 4.4.
We say that $X\subset\mathbb{R}^{n}$ is a Lipschitz graph at infinity if there is a compact subset $K$ such that $X\setminus K$ is the graph of a Lipschitz mapping $\phi\colon U\to\mathbb{R}^{n-k}$, where $U$ is an open subset of $\mathbb{R}^{k}$.
The next result was essentially proved by Birbrair, Fernandes and Jelonek [3], but they do not use the terminology of the fifth tangent cone at infinity. For completeness, we present a proof here.
Proposition 4.5.
Let $X\subset\mathbb{R}^{n}$ be a pure $d$-dimensional closed definable set in an o-minimal structure. If $\dim C_{5}(X,\infty)=d$, then $X$ is a Lipschitz graph at infinity on $\pi(X\setminus K)$.
Proof.
Let $P$ be a $d$-dimensional linear subspace such that the orthogonal projection $\pi\colon\mathbb{R}^{n}\to P$ satisfies $\pi^{-1}(0)\cap C_{5}(X,\infty)=\{0\}$.
Claim 4.5.1.
There are $C>0$ and a compact subset $K$ such that $\|\pi(x)-\pi(y)\|\geq C\|x-y\|$ for all $x,y\in X\setminus K$.
Proof of Claim 4.5.1.
Since $\pi$ is a definable mapping, if the claim was not true, then there would be a pair of definable arcs $\gamma_{1},\gamma_{2}\colon[r,+\infty)\to X$ such that $\lim\limits_{t\to+\infty}\|\gamma_{1}(t)\|=\lim\limits_{t\to+\infty}\|\gamma_{2}(t)\|=+\infty$ and
$\lim\limits_{t\to+\infty}\frac{\|\varphi(\gamma_{1}(t))-\varphi(\gamma_{2}(t))\|}{\|\gamma_{1}(t)-\gamma_{2}(t)\|}=0.$
For $v=\lim\limits_{t\to+\infty}\frac{\gamma_{1}(t)-\gamma_{2}(t)}{\|\gamma_{1}(t)-\gamma_{2}(t)\|}$, we have that $v\in\pi^{-1}(0)\cap C_{5}(X,\infty)$, which is a contradiction, since $\pi^{-1}(0)\cap C_{5}(X,\infty)=\{0\}$ and $v\not=0$. ∎
Thus, we can define a function $f:\pi(X\setminus K)\to\mathbb{R}^{n-k}$, such that $\Gamma(f)=X\setminus K$, with $f$ Lipschitz. If $f$ were not Lipschitz for any compact $K$, then for every given $k\in\mathbb{N}$, there exist $(x_{k}),(y_{k})\subset\pi(X\setminus B_{k}(0))$ such that
${\|f(x_{k})-f(y_{k})\|\over\|(x_{k}-y_{k})\|}>k,$
but since $((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\over\|((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\|$ has norm 1, it converges up to a subsequence. But
${x_{k}-y_{k}\over\|((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\|}\leq{x_{k}-y_{k}\over\|(f(x_{k})-f(y_{k}))\|}\to 0,$
then the sequence $((x_{k},f(x_{k}))-(y_{k},f(y_{k})))\over|((x_{k},f(x_{k}))-(y_{k},f(y_{k})))|$ converges to a vector that lies in $\pi^{-1}(0)\cap dimC_{5}(X,\infty)=\{0\}$, but on the other hand has norm $1$, a contradiction. Therefore, $f$ is Lipschitz. ∎
The converse of the above proposition does not hold.
Example 4.6.
Let $X=\{(x,y)\in R^{2};y=|x|\}$. We have that $X$ is the graph of a Lipschitz function, but $C_{5}(X,\infty)=\{(x,y)\in R^{2};-|x|\leq y\leq|x|\}$.
As a consequence, we obtain one of the main theorems of [6], which also follows from 4.5 and [8, Theorem 3.8] (see also [18, Theorem 1.6]):
Corollary 4.7.
Let $X\subset\mathbb{C}^{n}$ be a pure $d$-dimensional complex algebraic set. If $\dim C_{5}(X,\infty)=2d$, then $X$ is an affine linear subspace.